OpenStudy (anonymous):

a number added to itself seven additional times is fourteen more than the number. Whats the number?

OpenStudy (kobeni-chan):

2

OpenStudy (swissgirl):

x+7x=14+x solve for x

OpenStudy (anonymous):

2

OpenStudy (anonymous):

x+7x= 14+x 7x+x is 8x=14+x 8x-x is now 7x=14 lastly 14/7 is 2 x=2